If you were going to make a ladder to climb a top the Triceratops back, Providing hthey would let you and the distance from the Triceratops to where the ladder will go to the sand is 24ft. How long would the ladder be.
Triceratops is 26 ft long and 10 ft high.

Answer by checkley71(8403)

FIRST YOU NEED TO DRAW A DIAGRAM OF THE KNOWN FACTS. THE DISTANCE FROM THE TRICERATOPS IS 24 FEET & THE HEIGHT IS 10 FEET. THIS GIVES YOU A RIGHT TRIANGLE FOR WHICH YOU NEED TO FINE THE HYPOTENUSE.
-------------------------------------------------------------------------
USING THE PYTHAGOREAN FORMULA A^2+B^2=C^2
WE HAVE 10^2+24^2=C^2
100+576=C^2
C^2=676
C=SQRT676
C=26 FEET IS THE LENGTH OF THE LADER REQUIRED TO GET TO THE TOP.